OpenStudy (anonymous):

solve the equation. 1+1 over x = 30 over x^2

OpenStudy (anonymous):

hello :)

OpenStudy (s3a):

hi

OpenStudy (s3a):

(1+1)/30/90

OpenStudy (s3a):

= 2/30 * 1/90 = 2/(30*90)= 2/2700

OpenStudy (anonymous):

?

OpenStudy (anonymous):

nswers are

OpenStudy (s3a):

Maybe i misunderstood the question.

OpenStudy (s3a):

And maybe you should of put the answers here in the first place lol ;)

OpenStudy (anonymous):

-6,5 -5,6 5,6 -1/6,1/5

OpenStudy (anonymous):

haha i wanted to make you work for it

OpenStudy (s3a):

lol

OpenStudy (s3a):

is there a way you can write this by hand? because that's probably why i didn't understand you

OpenStudy (s3a):

or using a mouse

OpenStudy (anonymous):

yes one second

OpenStudy (s3a):

ight

OpenStudy (anonymous):